OpenStudy (anonymous):

Let f(x) = 9/x The slope of the tangent line to the graph of f(x) at the point (-2, -9/2) is -2.25. The equation of the tangent line to the graph of f(x) at (-2, -9/2) is y=mx+b for m = -2.25 b = ?

OpenStudy (amistre64):

y = mx -mPx +Py y = -2.25x +(2.25)(-2) -(9/2)

OpenStudy (amistre64):

b = -4.5 - 9/2 = -4.5 -4.5 = -9

OpenStudy (anonymous):

ha thanks again, man. What's the universal formula for getting the intercept? I can figure everything else out but that :/

OpenStudy (amistre64):

well; in my universe; I use -m(Px) + Py; where; m=slope and P(x,y)

OpenStudy (amistre64):

if you algebrate the point slope form you get this thing y - Py = m(x - Px) y = mx -mPx + Py; where b = -mPx +Py

OpenStudy (anonymous):

ooohh, I was messing up on the symbols -_- got it! cheers :]